OpenStudy (anonymous):

Write an equation of the line parallel to y=2/3x+3 and passes through the point (9,-10).

OpenStudy (mathstudent55):

\(y = \frac{2}{3}x + 3\) is in slope-intercept form, \(y = mx + b\) where m = slope and b = y-intercept. What is the slope of a parallel line?

OpenStudy (anonymous):

y+10=2/3(x-9) 3y+30=2(x-9) 3y+30=2x-18 3y=2x-48 y=2/3-16